# version 1.0

### v6

@@ 10.2.2 Closed-Loop Optimal Predictors
Let the control sequence $\mathbf{u}_{t}$ be an arbitrary affine function of $\mathbf{y}_{1:t}$. There exists an efficiently computable sequence of linear predictors $\hat{\mathbf{x}}_{t+1}$ defined recursively as
@@
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
@@ 10.2.3 Tools from Linear Regression
In the remaining proof, we use this observation to construct an optimal least-squares estimator $\hat{\mathbf{x}}_{t+1}$ for $\mathbf{x}_{t+1}$ given $\hat{\mathbf{x}}_{t}$. In particular, we know by induction hypothesis that
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L\in\prod^{L}}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2}.$
But, is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 10.6).
@@ 10.2.4 Analysis
We are now ready to prove the main property of the Kalman filter below.
@@
$\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}.$
Since for any $s<t$
@@
$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L\in\Pi^{L}}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2}.$
But, when predicting $\mathbf{x}_{t+1}$, we have more information. In particular, we also have $\mathbf{y}_{t}$. Consider random variables $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$. Can we use this covariate alone to predict $\mathbf{x}_{t+1}$? Note that
$\mathbb{E}\left[(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]$ $=\Sigma_{v}+C\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]C^{\top}$
$=\Sigma_{v}+C\Sigma_{t}C^{\top},$
@@
In the preceding paragraphs, we have constructed two least-squares predictors for $\hat{\mathbf{x}}_{t+1}$. Observe that for any $s<t$
@@
$\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{w}_{t}+L_{t}\mathbf{v}_{t}.$
Finally, we give a recursive relation on $\Sigma_{t}$ as follows–
@@
$=A\Sigma_{t}A^{\top}+\Sigma_{x}+L_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})L_{t}^{\top}-2A\Sigma_{t}C^{\top}L_{t}^{\top}$
$=A\Sigma_{t}A^{\top}-A\Sigma C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma A^{\top}+\Sigma_{x}.$
@@ 10.2.5 Infinite Horizon Kalman Filtering
In the infinite-horizon setting, the steady-state Kalman filter can be described by the following linear dynamical system
@@
where $L=A\Sigma(C\Sigma C^{\top}+\Sigma_{y})^{-1}$ and $\Sigma$ satisfies the Riccati equation
@@
However, when does a solution to this Riccati equation exist? A sufficient condition for this is the observability of the underlying dynamical system. In particular, any observable partially-observed linear dynamical system admits a solution to the Riccati equation listed above, and hence, has a steady-state Kalman filter. Mirroring the link between controllability and stabilizability, observability implies $\rho(A-LC)<1$.
10.2.6 Linear Predictors and the Kalman Filter
For observable linear dynamical systems, the Kalman filter is contained in the class of linear predictors. This is formally stated in the following lemma, whose proof is left as Exercise 10.6.
@@ 10.2.6 Linear Predictors and the Kalman Filter
Consider any partially-observed linear dynamical system $(A,B,C)$ where the Kalman filtering matrix $L$ is such that the system ALCA-LC is $\gamma$-stable. The Kalman filtering policy is $\varepsilon$-contained in the class of linear policies $\Pi^{L}_{h,k,\kappa}$, where $h=k=\gamma\log\frac{\gamma T}{\varepsilon}$ and $\kappa=\gamma^{3}$.
@@ 10.3 Bayes-Optimality under Gaussian Noise
In this subsection, we further restrict $\mathbf{w}_{t},\mathbf{v}_{t}$ to be i.i.d. draws from Gaussian distributions, and proceed to establish that the Kalman filter in such cases enjoys a stronger notion of optimality; it produces the best estimator (under a least-squares loss) of the state given past observations among the class of all (not just linear) predictors.  [^1] This is also called the Bayes-optimal predictor. See bibliographic materials for more details and literature. . The setting we consider is that of Gaussian perturbations $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ and known systems matrices $(A,B,C)$. In particular, we consider the case where $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d. Gaussian and $x_{0}=0$. Given the realizations $\mathbf{y}_{1},\mathbf{y}_{2}\dots,\mathbf{y}_{t}$ under some control input sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots\mathbf{u}_{t}$, we would like to compute a prediction $\hat{\mathbf{y}}_{t+1}$ that minimizes the mean square error, which is simply the expected observation as per
$\hat{\mathbf{y}}_{t+1}=\mathbb{E}[\mathbf{y}_{t+1}|\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
Since $\hat{\mathbf{y}}_{t+1}=C\hat{\mathbf{x}}_{t+1}$, we will recursively compute $\hat{\mathbf{x}}_{t+1}$ where
$\hat{\mathbf{x}}_{t+1}=\mathbb{E}[\mathbf{x}_{t+1}|\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
@@
for a sequence of matrices $L_{t}\in{\mathbb{R}}^{d_{x}\times d_{y}}$ that can be efficiently computed given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
@@
$\mathbf{x}_{t}|(\mathbf{y}_{1:t},\mathbf{u}_{1:t-1})$ $\sim\mathcal{N}(\hat{\mathbf{x}}_{t}+\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),\Sigma^{\prime}_{t}),$
@@
$=(A-\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}\mathbf{y}_{t}.$
@@ 10.4 Conclusion
In this chapter, we explored the Kalman filter as a fundamental tool for state estimation in dynamical systems, particularly in the presence of noise and uncertainty. We established its optimality in the least-squares sense and examined its connections to online convex optimization, reinforcing its relevance in both classical control and modern learning-based paradigms.
Key insights from this chapter include the role of observability in ensuring accurate state estimation, the derivation of the Kalman filter update equations, and the recursive structure that allows for efficient real-time implementation. We also presented the filter’s steady-state behavior, particularly in the infinite-horizon setting, where it converges to a solution governed by the Riccati equation.
Beyond its theoretical guarantees, the Kalman filter is widely used in practical applications ranging from autonomous systems and robotics to finance and neuroscience. However, challenges remain, particularly in scenarios with non-Gaussian noise, time-varying dynamics, or adversarial disturbances. These limitations motivate extensions such as robust filtering, nonlinear filtering (e.g., the extended and unscented Kalman filters) and learning-based adaptations that integrate statistical inference with modern machine learning techniques.
The next chapter will transition from the classical filtering framework to spectral filtering, providing an alternative approach to state estimation that leverages spectral methods for efficient learning in dynamical systems.
@@ 10.5 Bibliographic Remarks
In 1960, Rudolph Kalman gave a state-space solution to the filtering problem that was highly amenable to further extensions [Kal60]. See [AM91] for a historical perspective on its development and aftermaths.
For unknown systems, in the stochastic case, one may estimate the system matrices, as we have discussed before, before applying optimal filtering. Improper learning approaches to this problem were studied in [KMTM19].
@@
  • Nonlinear Filtering Extensions. The Kalman filter assumes that the dynamics is linear system and the noise is Gaussian, which limits its effectiveness in nonlinear settings. The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, linearization can introduce significant errors in highly nonlinear systems. Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
  • Particle Filtering (Sequential Monte Carlo Methods)
    For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
  • Adversarial & Robust Filtering
    Robust filtering methods seek to provide guarantees under adversarial noise. The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-squared error, making it suitable for applications with unknown disturbances.
Although the Kalman filter remains a cornerstone of state estimation, these alternative methods provide valuable extensions and improvements for nonlinear, adversarial, and data-driven settings. In the next chapter, we will study an alternative using a technique which circumvents the need for state estimation altogether and has stronger guarantees for sequential prediction.
@@ 10.6 Exercises
10.1. In this exercise we prove Theorem 10.2 in two parts:
@@
    Show that if the rank of the Kalman observability matrix is less than $d_{\mathbf{x}}$, then there are two initial states that after $d_{\mathbf{x}}$ iterations map to the same state. Hence, the initial state is not uniquely determined.
@@
    In contrast, show that if the rank of $\hat{K}$ is full, then at any time $t$, $\mathbf{x}_{t}$ determines $\mathbf{x}_{0}$ uniquely.
@@
10.4. In this exercise we conclude that the Kalman filter can be improperly learned by the class of linear predictors. Prove this fact as formally stated in Lemma 10.6. Conclude that the Kalman filter can be learned using Algorithm 6.
@@
In the previous chapter, we explored two major approaches for state estimation and prediction in unknown linear dynamical systems:
@@
    Online learning of linear predictors, which offers a relaxation of the Kalman filtering approach by directly learning predictive models without explicit system identification. Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs. Although this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the stability constant of the system, which can be large and unknown, making learning impractical for unstable or poorly conditioned systems.

### v8

@@ 10.2 The Kalman Filter
Let the control sequence $\mathbf{u}_{t}$ be an arbitrary affine function of $\mathbf{y}_{1:t}$. There exists an efficiently computable sequence of linear state estimates $\hat{\mathbf{x}}_{t+1}$ defined recursively as
@@
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L}\mathbb{E}\|\mathbf{x}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2},$
where the minimum ranges over all linear maps from $(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$ to $\mathbb{R}^{d_{\mathbf{x}}}$.
@@ 10.2.1 Tools from Linear Regression
In the remaining proof, we use this observation to construct an optimal least-squares estimator $\hat{\mathbf{x}}_{t+1}$ for $\mathbf{x}_{t+1}$ from the available information. In particular, by the induction hypothesis,
$\mathbb{E}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}\leq\min_{L}\mathbb{E}\|\mathbf{x}_{t}-L\left[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}\right]\|^{2},$
where the minimum ranges over all linear maps from $(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$ to $\mathbb{R}^{d_{\mathbf{x}}}$. But is $\hat{\mathbf{x}}_{t}$ any good for predicting $\mathbf{x}_{t+1}$? To answer this, we note the following structural characterization of least-squares (Exercise 10.6).
@@ 10.2.2 Analysis
We are now ready to prove the main property of the Kalman filter below. Throughout this section, assume that $\Sigma_{y}\succ 0$.
@@
$\Sigma_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right].$
Since for any $s<t$,
@@
$\mathbb{E}\|\mathbf{x}_{t+1}-A\hat{\mathbf{x}}_{t}-B\mathbf{u}_{t}\|^{2}\leq\min_{L}\mathbb{E}\|\mathbf{x}_{t+1}-B\mathbf{u}_{t}-L[\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1}]\|^{2},$
where the minimum ranges over all linear maps from $(\mathbf{y}_{1:t-1},\mathbf{u}_{1:t-1})$ to $\mathbb{R}^{d_{\mathbf{x}}}$.
But, when predicting $\mathbf{x}_{t+1}$, we have more information. In particular, we also have $\mathbf{y}_{t}$. Consider the random variable $\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}$. Can we use this covariate alone to predict $\mathbf{x}_{t+1}$? Note that
$\mathbb{E}\left[(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t})^{\top}\right]$ $=\Sigma_{y}+C\mathbb{E}\left[(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})^{\top}\right]C^{\top}$
$=\Sigma_{y}+C\Sigma_{t}C^{\top},$
@@
In the preceding paragraphs, we have constructed two least-squares predictors for ${\mathbf{x}}_{t+1}$. Observe that for any $s<t$,
@@
$\mathbf{x}_{t+1}-\hat{\mathbf{x}}_{t+1}=(A-L_{t}C)(\mathbf{x}_{t}-\hat{\mathbf{x}}_{t})+\mathbf{w}_{t}-L_{t}\mathbf{v}_{t}.$
Finally, we give a recursive relation on $\Sigma_{t}$ as follows:
@@
$=A\Sigma_{t}A^{\top}+\Sigma_{x}+L_{t}(C\Sigma_{t}C^{\top}+\Sigma_{y})L_{t}^{\top}-A\Sigma_{t}C^{\top}L_{t}^{\top}-L_{t}C\Sigma_{t}A^{\top}$
$=A\Sigma_{t}A^{\top}-A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}C\Sigma_{t}A^{\top}+\Sigma_{x}.$
@@ 10.2.3 Infinite Horizon Kalman Filtering
In the infinite-horizon setting, when the covariance recursion converges, the Kalman filter admits a steady-state description with a fixed gain matrix $L$. The resulting estimator is itself a linear dynamical system driven by the control inputs and the incoming observations:
@@
where $L=A\Sigma C^{\top}(C\Sigma C^{\top}+\Sigma_{y})^{-1}$ and $\Sigma$ satisfies the Riccati equation
@@
The existence of a stabilizing solution to this Riccati equation is a classical question. Under standard assumptions—typically detectability of $(A,C)$ together with an appropriate stabilizability condition for the state-noise model—the Riccati equation admits a stabilizing solution. In that case the steady-state Kalman gain $L$ is well defined and the closed-loop matrix ALCA-LC is stable. We do not pursue the sharp conditions here.
10.2.4 Linear Predictors Approximate the Kalman Filter
At first sight, the steady-state Kalman filter appears to lie outside the class of predictors studied in Chapter 9 because it maintains an internal state estimate and uses feedback from the newest observation. The next lemma gives a relationship between the class of predictors. Once the closed-loop matrix ALCA-LC is stable, the Kalman recursion can be unrolled and truncated, yielding a finite-memory linear predictor with exponentially small approximation error. This is the closed-loop analogue of the truncation phenomenon from Chapter 9.
@@ 10.2.4 Linear Predictors Approximate the Kalman Filter
Fix $M\geq 1$. Consider the family of steady-state Kalman filters
$\hat{\mathbf{x}}_{t+1}=F\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+L\mathbf{y}_{t},\qquad\hat{\mathbf{y}}_{t}=C\hat{\mathbf{x}}_{t},$
with $\hat{\mathbf{x}}_{1}=0$, where F=ALCF=A-LC satisfies
$\|F^{s}\|\leq Me^{-s/M}\qquad\text{for all }s\geq 0,$
and $\|B\|+\|L\|+\|C\|\leq M$, as wel as $\max_{t\in[T]}\|\mathbf{u}_{t}\|\leq M,\qquad\max_{t\in[T]}\|\mathbf{y}_{t}\|\leq M$. Then the class $\Pi^{L}_{h,h,\kappa}$ $\varepsilon$-approximates this family over horizon $T$ provided
$h\geq M\log\!\left(\frac{4M^{5}}{\varepsilon}\right),\qquad\kappa\geq 4M^{4}.$
@@ 10.3 Bayes-Optimality under Gaussian Noise
The previous section established the Kalman recursion as the optimal linear closed-loop estimator. Under Gaussian noise, this linear optimality strengthens to full Bayes optimality. [^1] This is also called the Bayes-optimal predictor. See bibliographic materials for more details and literature. In this section we assume that $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are i.i.d. Gaussian and show that the Kalman filter computes the conditional expectation of the hidden state, and hence the mean-square optimal predictor of the next observation.
Concretely, assume that $\mathbf{w}_{t}\sim\mathcal{N}(0,\Sigma_{x})$ and $\mathbf{v}_{t}\sim\mathcal{N}(0,\Sigma_{y})$ are i.i.d. Gaussian and that $\mathbf{x}_{0}=0$. Given realizations $\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{t}$ under a control sequence $\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{t}$, we would like to compute
$\hat{\mathbf{y}}_{t+1}=\mathbb{E}[\mathbf{y}_{t+1}\mid\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
Since $\hat{\mathbf{y}}_{t+1}=C\hat{\mathbf{x}}_{t+1}$, it suffices to recursively compute
$\hat{\mathbf{x}}_{t+1}=\mathbb{E}[\mathbf{x}_{t+1}\mid\mathbf{y}_{1:t},\mathbf{u}_{1:t}].$
@@
for a sequence of matrices $L_{t}\in\mathbb{R}^{d_{\mathbf{x}}\times d_{y}}$ that can be efficiently computed given $(A,B,C)$, and is independent of the realizations $\mathbf{u}_{1:T},\mathbf{y}_{1:T}$.
@@
$\mathbf{x}_{t}|(\mathbf{y}_{1:t},\mathbf{u}_{1:t-1})$ $\sim\mathcal{N}\Big(\hat{\mathbf{x}}_{t}+\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}(\mathbf{y}_{t}-C\hat{\mathbf{x}}_{t}),\Sigma^{\prime}_{t}\Big),$
@@
$=(A-\underbrace{A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}C)\hat{\mathbf{x}}_{t}+B\mathbf{u}_{t}+\underbrace{A\Sigma_{t}C^{\top}(C\Sigma_{t}C^{\top}+\Sigma_{y})^{-1}}_{\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}L_{t}}\mathbf{y}_{t}.$
@@ 10.4 Conclusion
This chapter extends the story of Chapter 9 from open-loop prediction to closed-loop state estimation. The main conceptual point is that feedback changes the benchmark but not the convex relaxation: the Kalman filter is the natural optimal closed-loop predictor, yet stable Kalman filters are still approximable by finite-memory linear predictors.
We first introduced observability as the structural condition that makes state estimation possible. We then derived the Kalman recursion and its steady-state form, and showed that under Gaussian noise it is Bayes-optimal. Finally, we connected this classical estimator back to the learning perspective of the previous chapter by showing that finite-memory linear predictors can approximate stable Kalman filters. Consequently, the online learning machinery developed in Chapter 9 continues to apply in partially observed stochastic systems.
The next chapter moves from Kalman filtering to spectral filtering. There the emphasis shifts again from explicit state estimation to improper learning of predictors, but the central theme remains the same: exploit linear structure while avoiding direct nonconvex identification of the underlying dynamical system.
@@ 10.5 Bibliographic Remarks
In 1960, Rudolph Kalman gave a state-space solution to the filtering problem that was highly amenable to further extensions [KAL60]. See [AM91] for a historical perspective on its development and aftermath.
For unknown systems, in the stochastic case, one may estimate the system matrices, as we have discussed before, before applying optimal filtering. Improper learning approaches to this problem were studied in [KMT+19].
@@
  • Nonlinear Filtering Extensions. The Kalman filter assumes that the dynamics is linear and the noise is Gaussian, which limits its effectiveness in nonlinear settings. The Extended Kalman Filter (see, for example, [UA21]) addresses this issue by linearizing the dynamics at each time step via a first-order Taylor series expansion. However, linearization can introduce significant errors in highly nonlinear systems. Notable extensions using more sophisticated sampling include the Unscented Kalman Filter [JU04], leading to better estimation accuracy in strongly nonlinear systems.
  • Particle Filtering (Sequential Monte Carlo Methods). For highly nonlinear and non-Gaussian settings, particle filters [GSS93] provide a flexible alternative. These methods approximate the posterior distribution using a set of weighted samples (particles) and employ importance sampling and resampling techniques.
  • Adversarial & Robust Filtering. Robust filtering methods seek to provide guarantees under adversarial noise. The $H_{\infty}$ filter [BB08] minimizes the worst-case estimation error rather than the mean-squared error, making it suitable for applications with unknown disturbances.
In the next chapter, we will study an alternative using a technique which circumvents the need for state estimation altogether and has dimension-free guarantees for certain cases of sequential prediction.
@@ 10.6 Exercises
10.1. In this exercise we prove Theorem 10.2.
@@
    Suppose $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))<d_{\mathbf{x}}$. Show that there exists a nonzero vector $v\in\ker(\hat{K}(A,C))$. Conclude that the two initial states $\mathbf{x}_{0}$ and $\mathbf{x}_{0}+v$ generate the same observation sequence under the same input sequence, and hence the system is not observable.
@@
    Suppose $\mathop{\mbox{\rm rank}}(\hat{K}(A,C))=d_{\mathbf{x}}$. Show that the map
    $\mathbf{x}_{0}\mapsto\begin{bmatrix}C\mathbf{x}_{0}\\ CA\mathbf{x}_{0}\\ \vdots\\ CA^{d_{\mathbf{x}}-1}\mathbf{x}_{0}\end{bmatrix}$
    is injective. Conclude that, given the input sequence, the first $d_{\mathbf{x}}$ observations uniquely determine the initial state.
@@
10.4. Prove Lemma 10.6 by unrolling the steady-state Kalman recursion and truncating the resulting linear predictor. Conclude that, under the assumptions of the lemma, the family of stable steady-state Kalman filters can be learned improperly using Algorithm 6.
@@
In the previous chapters, we explored two major approaches for state estimation and prediction in unknown linear dynamical systems:
@@
    Online learning of linear predictors, which offers a relaxation of the Kalman filtering approach by directly learning predictive models without explicit system identification. Instead of estimating the full state and transition matrices, this approach constructs a forecasting model using past inputs and outputs. Although this avoids the complexity of full system identification, it introduces a major limitation: the number of parameters required depends on the hidden dimension of the system, or the stability constant of the system, which can be large and unknown, making learning impractical for high-dimensional or poorly conditioned systems.